OpenStudy (ittybitty14):

FAN & MEDAL!!!!!!!! Arrange these numbers least from greatest: 3 / 7, -5 /6, - 3 / 5, -7 / 9

OpenStudy (ghostgate):

Okay, you have to figure out there decimal forms, a calculator would be very useful here. 3/7 = 0.4285_ -5/6 = -0.8333_ -3/5 = -0.6 -7/9 = -0.7777_ Knowing this, what do you think the order should be?

OpenStudy (ittybitty14):

-3/5, 3/7,-7/9,-5/6?

OpenStudy (ghostgate):

Okay negative numbers are lower than positive numbers, so: -5/6 is the lowest number and should be your starting number. So, no that's not the order.

OpenStudy (ghostgate):

Think of it this way, negative numbers are from the left and positive numbers are from the right. You have:

OpenStudy (shansha1):

the negative numbers are lower than positive numbers, right so: -5/6 is the lowest number and should be your starting number. So, no that's not the order.

OpenStudy (ghostgate):

Lol - That's basically what I said.

OpenStudy (shansha1):

ight

OpenStudy (ittybitty14):

Thank you for your help, @Ghostgate I figured it out and got it right.

OpenStudy (ghostgate):

You're welcome. ;)